# ReviewGraph: A Knowledge Graph Embedding Based Framework for Review Rating Prediction with Sentiment Features

### v1

@@ Abstract
In the hospitality industry, understanding the factors that drive customer review ratings is critical for improving guest satisfaction and business performance. This work proposes ReviewGraph for Review Rating Prediction (RRP), a novel framework that transforms textual customer reviews into knowledge graphs by extracting subject–predicate–object triples and associating sentiment scores. Using graph embeddings (Node2Vec) and sentiment features, the framework predicts review rating scores through machine learning classifiers. We compare ReviewGraph performance with traditional NLP baselines (such as Bag of Words, TF-IDF, and Word2Vec) and large language models (LLMs), evaluating them in the HotelRec dataset. In comparison to the state of the art literature, our proposed model performs similar to their best performing model but with lower computational cost (without ensemble). While ReviewGraph achieves comparable predictive performance to LLMs and outperforms baselines on agreement-based metrics such as Cohen’s Kappa, it offers additional advantages in interpretability, visual exploration, and potential integration into Retrieval-Augmented Generation (RAG) systems. This work highlights the potential of graph-based representations for enhancing review analytics and lays the groundwork for future research integrating advanced graph neural networks and fine-tuned LLM-based extraction methods. We will share ReviewGraph output and platform open-sourced at https://github.com/aaronlifenghan/ReviewGraph
@@ I Introduction
In the past, consumers had a limited way to measure the quality of services within the hospitality industry. The main external source of information for them in this regard were professional reviewers. As the internet age went on, however, professional reviewers are increasingly replaced by other consumer reviewers for this purpose [1]. Consumer reviews affect up to 50 percent of booking decisions [2]. It also affects a multitude of metrics regarding hotel performance, among which are: occupancy rates, revenue per available room, guest satisfaction, and brand reputation. In particular, negative reviews affect these metrics strongly [3]. Taking these facts into account, it is of great interest for hotels to be able to accurately estimate what factors affect their review ratings.
We propose a new framework called ReviewGraph that will help hoteliers, and business owners generally by converting their reviews into a knowledge graph. We believe that this will allow a more context-aware and granular representation of what reviews say, allow users to inspect what causes low-rated reviews more easily, and in the future could be used for Retrieval Augmented Generation of user reviews to quickly have LLMs answer questions about their reviews. To investigate this topic, we have formulated the following research questions: Research Question 1: How can knowledge graph embeddings and sentiment prediction models contribute to customer review based rating predictions? Research Question 2: How do current LLMs perform on review score prediction tasks, in comparison to task-specific NLP models?
In summary, the main contributions of this paper are: One, the design and implementation of a proof of concept of the ReviewGraph framework. Two, an empirical evaluation comparing our implemented framework to classical baselines and an LLM baseline. And three, an analysis of the trade-offs between these baselines and our proposed model.
@@ II-A1 Word Embeddings and Language Models
Traditional Natural Language Processing has mainly relied on embedding techniques such as Bag-of-Words or TF-IDF, which ignored word order or semantic similarity. The introduction of newer embedding techniques such as Word2Vec [4] addressed this limitation by representing words as vectors where semantic similarity is reflected by its proximity. Word2Vec paved the way for embedding techniques and language models that also consider the order within the sentence dynamically such as ELMo [5] and BERT [6].
@@ II-A2 Large Language Models
More recently, Large Language Models (LLMs) such as GPT-4 [7] are being used for a wide variety of tasks which have been trained on massive datasets, and transformer-based architectures. Because of their architecture and large dataset, they are increasingly used not just for generating text, but also for extraction tasks. In this study, we will also utilize LLMs to predict review ratings.
@@ II-A3 Classification task in NLP
Text classification is a common problem in the field of NLP defined as the use of a machine learning model, referred to as a classifier, that is trained and used to differentiate between predefined classes based on features that have been extracted from the text in some manner [8]. The accuracy of the task depends on the granularity of the method, and on how clearly separable the different text units are from one another [8].
@@ II-A4 Knowledge Graphs
Knowledge graphs are structured representations of information in the form of a graph. They use a graph-based data model to capture knowledge at scale using information from many different sources. It allows the maintainers to forego a schema and let the database evolve organically [9]. Knowledge Graphs have garnered extra attention in recent times because of its potential for improving RAG (Retrieval Augmented Generation) using the GraphRAG approach [10].
@@ II-B1 Customer Review Predictions
Kumar et al. (2024) present an ensemble learning approach for predicting hotel ratings based on user-generated review text content [11]. The review data they used is from the same dataset that this paper used. The research makes use of classical word embedding methods such as Bag of Words, TF-IDF, and Word2Vec, and compares these embedding methods to each other. To predict the ratings they make use of an ensemble of classifier models, among which are: Stochastic Gradient Descent Classifier, Logistic Regression, Support Vector Classifier, Random Forest Classifier, K-Nearest Neighbors, and Decision Tree. For the ensemble method, Kumar et al. used hard voting combined with a majority vote to decide the predicted rating score. The only metric they used is accuracy score, and their final highest accuracy score was 57% with Bag of Words embedding plus the ensemble classifier model. In our investigation, we aimed for this accuracy score, and we used their methods as baselines to compare our results to using our new method. As well as using the same dataset. Additionally, they did not make use of sentiment scores while we do.
@@ II-B2 Review Visualisation in Hospitality
As more users generate online reviews, it becomes increasingly difficult for both consumers and businesses to quickly summarize and understand opinions about products and services. One of the earlier works in this space proposed an interactive visualization system that presents commonly mentioned topics and their associated average sentiments using aspect-based sentiment analysis and topic modeling [12].
More recently, Huth et al. introduced ViSCitR (Visual Comparison of Cities through Reviews), a system designed to improve the analysis of review data by providing summaries of both positive and negative opinions for each topic [13]. ViSCitR combines several coordinated views. This includes a geographic map, numeric rating charts across fixed categories, temporal rating trends, and textual summaries of frequently mentioned points. However, ViSCitR also has key limitations. Its design emphasizes only recurring patterns, deliberately discarding less frequent or minority opinions to maintain clarity, and it reduces sentiment to positive/negative counts without modeling the intensity of this senetiment. Additionally, all review aspects must fit into a predefined schema of topics, meaning more niche topics may be generalized into generic topics. For example, the unclean bathroom may be generalized as cleanliness (or bathroom). The tool also does not offer predictive analytics or structural analysis to uncover hidden patterns or influential factors beyond the presented summaries.
@@ II-B3 Triple Extraction for Knowledge Graph Construction
To be able to generate the embeddings for our model, we first need to construct the knowledge graph from the textual review data. Most research in this space is focused on NER (Named Entity Recognition) and relationship extraction [14]. As we are not just trying to extract commonly mentioned entities (though that is an important part of it), these methods are not satisfactory. It would not recognise sentences such as "it was great" or "it was terrible", missing important context.
As an alternative we investigated Large Language Model based method such as the one by Zhang et. al. (2024) [15]. This method involves fine-tuning an LLM to extract full "triples" from text data as follows: ⟨ subject, predicate, object ⟩. These methods, while promising, are computationally intensive and time-consuming. Moreover, our study already investigates the effectiveness of LLMs in review rating prediction, making the additional complexity of fine-tuning less justifiable.
Given these considerations, we opted to use the Stanford CoreNLP Open Information Extraction (OpenIE) method [16]. It is a non-LLM method that extracts full triples as ⟨ subject, predicate, object ⟩ similarly to the fine-tuned LLM method from the previous paragraph.
@@ II-B4 Knowledge Graph Usage in Hospitality and Business
In the hospitality industry most investigations on the application of knowledge graphs has involved improving recommendation systems to better recommend hotels to users [17]. In the broader business context, Knowledge Graphs have been attracting increasing attention because of their inherent usefulness to aid with Retrieval Augmented Generation [10].
@@ III-A Dataset
The dataset we used is a fraction of the HotelRec [18] dataset. Which contains roughly 50 million reviews on hotels from TripAdvisor. We used only the first 10,000 reviews in this dataset for this study. Each review in the database contains the following features:
@@
Among these 10,000 reviews were 59 different unique hotels. The average rating of all these reviews is 4.162, with a standard deviation of 1.08. The average length of a review is 133 characters, while the median review length is 94. Any reviews longer than 1000 words were removed from this chart for readability purposes.
@@ III-B1 Preprocessing and Embedding
The preprocessing pipeline is implemented using standard tools from the nltk [19] and scikit-learn [20] libraries. The result is a set of clean, tokenized text inputs ready for feature extraction using the respective embedding techniques.
@@ III-B2 Sentiment Analysis
Optionally, VADER [21] sentiment analysis can be used to generate a sentiment score for the review, and use this as an additional feature when training the model.
@@ III-C1 Preprocessing and Triple Extraction
Each review is first translated to English, if necessary, using automatic language detection and the Google Translate APwe [22]. Next, contractions are expanded, for example: “can’t” $\rightarrow$ “cannot”, to standardize expressions. The text is then cleaned by removing HTML tags, hyphens, apostrophes, and similar punctuation that may interfere with parsing.
@@ III-C2 Sentiment Analysis
Optionally, sentiment analysis can now be performed on the list of triples using the VADER sentiment analyzer [21]. This returns a sentiment score that can range from -1 to 1. This score is stored in the same csv file, in a separate column.
@@ III-C3 Graph Embedding
Once that is done, for each object and subject a node is created with a relationship created based on the predicate. Each of these nodes is connected to a node representing the review. The review node in turn is connected to a node representing the hotel. If the object/subject is in the list of amenities, it instead gets a special node type called "amenity", otherwise its node type is called "word". This is only relevant for the system visualization. Once this graph has been constructed using NetworkX [23]. It is then converted into a Neo4J [24] graph. This is to be able to visualize the graph with many more nodes than NetworkX supports, and to make use of its Node2Vec feature.
@@ III-D LLM-Based Model
Our LLM-based model works by asking the LLM to predict the ratings of a set of reviews, based on the initial "training set" it is given. So we first give it 200, or 2000 reviews with their corresponding ratings, and then based on those it has to predict the ratings for the other 9800, or 8000 reviews we give it next. To accomplish this task the LLM is allowed to write its own code.
@@ IV-B ReviewGraph Results
For the following experiments, we exclusively use the Node2Vec embedding method, as implemented in the Neo4j Graph Data Science Library [24]. The model was configured with the following parameters: 1 iteration, embedding dimension of 10, walk length of 80, return factor of 1, in-out factor of 1, and a window size of 10. The embeddings were generated for nodes with the label review, using relationships of type Any and orientation Any. No relationship weight property was applied in this setup. These values may change if mentioned.
@@ IV-E Discussion
V Conclusions
@@
V-A Limitations and Further Research
@@
Appendix A Sample of Extracted Triples
@@ Appendix B Key Implementation Code
Appendix C Complete Tables of Results
C-A Train Test Split Results
@@
C-B 10-Fold Cross Validation Results

### v2

@@ Abstract
In the hospitality industry, understanding the factors that drive customer review ratings is critical for improving guest satisfaction and business performance. This work proposes ReviewGraph for Review Rating Prediction (RRP), a novel framework that transforms textual customer reviews into knowledge graphs by extracting (subject, predicate, object) triples and associating sentiment scores. Using graph embeddings (Node2Vec) and sentiment features, the framework predicts review rating scores through machine learning classifiers. We compare ReviewGraph performance with traditional NLP baselines (such as Bag of Words, TF-IDF, and Word2Vec) and large language models (LLMs), evaluating them in the HotelRec dataset. In comparison to the state of the art literature, our proposed model performs similar to their best performing model but with lower computational cost (without ensemble). While ReviewGraph achieves comparable predictive performance to LLMs and outperforms baselines on agreement-based metrics such as Cohen’s Kappa, it offers additional advantages in interpretability, visual exploration, and potential integration into Retrieval-Augmented Generation (RAG) systems. This work highlights the potential of graph-based representations for enhancing review analytics and lays the groundwork for future research integrating advanced graph neural networks and fine-tuned LLM-based extraction methods. We will share ReviewGraph output and platform open-sourced on our GitHub page https://github.com/aaronlifenghan/ReviewGraph
@@ I Introduction
In the past, consumers had a limited way to measure the quality of services within the hospitality industry. The main external source of information for them in this regard were professional reviewers. As the internet age went on, however, professional reviewers are increasingly replaced by other consumer reviewers for this purpose [Viglia2016]. Consumer reviews affect up to 50 percent of booking decisions [Duan2016]. It also affects a multitude of metrics regarding hotel performance, among which are: occupancy rates, revenue per available room, guest satisfaction, and brand reputation. In particular, negative reviews affect these metrics strongly [gabbard2023]. Taking these facts into account, it is of great interest for hotels to be able to accurately estimate what factors affect their review ratings.
We propose a new framework called ReviewGraph  [^1] Peer-reviewed and Published version in ICKG-2025 proceedings https://cyprusconferences.org/ickg2025/ (The 16th IEEE International Conference on Knowledge Graphs, November 13-14, 2025, Limassol, Cyprus). that will help hoteliers, and business owners generally by converting their reviews into a knowledge graph. We believe that this will allow a more context-aware and granular representation of what reviews say, allow users to inspect what causes low-rated reviews more easily, and in the future could be used for Retrieval Augmented Generation of user reviews to quickly have LLMs answer questions about their reviews. To investigate this topic, we have formulated the following research questions: Research Question 1: How can knowledge graph embeddings and sentiment prediction models contribute to customer review based rating predictions? Research Question 2: How do current LLMs perform on review score prediction tasks, in comparison to task-specific NLP models?
In summary, the main contributions of this paper are: I), the design and implementation of a proof of concept of the ReviewGraph framework; II), an empirical evaluation comparing our implemented framework to classical baselines and an LLM baseline; and III), an analysis of the trade-offs between these baselines and our proposed model.
@@ II-A1 Word Embeddings and Language Models
Traditional Natural Language Processing has mainly relied on embedding techniques such as Bag-of-Words or TF-IDF, which ignored word order or semantic similarity. The introduction of newer embedding techniques such as Word2Vec [mikolov2013efficient] addressed this limitation by representing words as vectors where semantic similarity is reflected by its proximity. Word2Vec paved the way for embedding techniques and language models that also consider the order within the sentence dynamically such as ELMo [peters2018deep] and BERT [devlin2019bert].
@@ II-A2 Large Language Models
More recently, Large Language Models (LLMs) such as GPT-4 [openai2023gpt4] are being used for a wide variety of tasks which have been trained on massive datasets, and transformer-based architectures. Because of their architecture and large dataset, they are increasingly used not just for generating text, but also for extraction tasks. In this study, we will also utilize LLMs to predict review ratings.
@@ II-A3 Classification task in NLP
Text classification is a common problem in the field of NLP defined as the use of a machine learning model, referred to as a classifier, that is trained and used to differentiate between predefined classes based on features that have been extracted from the text in some manner [dogra2022complete] or prompting LLMs nowadays [cui2023medtem2]. The accuracy of the task depends on the granularity of the method, and on how clearly separable the different text units are from one another [dogra2022complete].
@@ II-A4 Knowledge Graphs
Knowledge graphs are structured representations of information in the form of a graph. They use a graph-based data model to capture knowledge at scale using information from many different sources. It allows the maintainers to forego a schema and let the database evolve organically [10.1145/3447772]. Knowledge Graphs have garnered extra attention in recent times because of its potential for improving RAG (Retrieval Augmented Generation) using the GraphRAG approach [edge2025localglobalgraphrag].
@@ II-B1 Customer Review Predictions
Kumar et al. (2024) present an ensemble learning approach for predicting hotel ratings based on user-generated review text content [Kumar2024]. The review data they used is from the same dataset that this paper used. The research makes use of classical word embedding methods such as Bag of Words, TF-IDF, and Word2Vec, and compares these embedding methods to each other. To predict the ratings they make use of an ensemble of classifier models, among which are: Stochastic Gradient Descent Classifier, Logistic Regression, Support Vector Classifier, Random Forest Classifier, K-Nearest Neighbors, and Decision Tree. For the ensemble method, Kumar et al. used hard voting combined with a majority vote to decide the predicted rating score. The only metric they used is accuracy score, and their final highest accuracy score was 57% with Bag of Words embedding plus the ensemble classifier model. In our investigation, we aimed for this accuracy score, and we used their methods as baselines to compare our results to using our new method. As well as using the same dataset. Additionally, they did not make use of sentiment scores while we do.
@@ II-B2 Review Visualisation in Hospitality
As more users generate online reviews, it becomes increasingly difficult for both consumers and businesses to quickly summarize and understand opinions about products and services. One of the earlier works in this space proposed an interactive visualization system that presents commonly mentioned topics and their associated average sentiments using aspect-based sentiment analysis and topic modeling [Ishizuka2015].
More recently, Huth et al. introduced ViSCitR (Visual Comparison of Cities through Reviews), a system designed to improve the analysis of review data by providing summaries of both positive and negative opinions for each topic [Huth2024]. ViSCitR combines several coordinated views. This includes a geographic map, numeric rating charts across fixed categories, temporal rating trends, and textual summaries of frequently mentioned points. However, ViSCitR also has key limitations. Its design emphasizes only recurring patterns, deliberately discarding less frequent or minority opinions to maintain clarity, and it reduces sentiment to positive/negative counts without modeling the intensity of this senetiment. Additionally, all review aspects must fit into a predefined schema of topics, meaning more niche topics may be generalized into generic topics. For example, the unclean bathroom may be generalized as cleanliness (or bathroom). The tool also does not offer predictive analytics or structural analysis to uncover hidden patterns or influential factors beyond the presented summaries.
@@ II-B3 Triple Extraction for Knowledge Graph Construction
To be able to generate the embeddings for our model, we first need to construct the knowledge graph from the textual review data. Most research in this space is focused on NER (Named Entity Recognition) and relationship extraction [ZHAO2023225]. As we are not just trying to extract commonly mentioned entities (though that is an important part of it), these methods are not satisfactory. It would not recognise sentences such as "it was great" or "it was terrible", missing important context.
As an alternative we investigated Large Language Model based method such as the one by Zhang et. al. (2024) [zhang-etal-2024-fine-tuning]. This method involves fine-tuning an LLM to extract full "triples" from text data as follows: ⟨ subject, predicate, object ⟩. These methods, while promising, are computationally intensive and time-consuming. Moreover, our study already investigates the effectiveness of LLMs in review rating prediction, making the additional complexity of fine-tuning less justifiable.
Given these considerations, we opted to use the Stanford CoreNLP Open Information Extraction (OpenIE) method [angeli2015leveraging]. It is a non-LLM method that extracts full triples as ⟨ subject, predicate, object ⟩ similarly to the fine-tuned LLM method from the previous paragraph.
@@ II-B4 Knowledge Graph Usage in Hospitality and Business
In the hospitality industry most investigations on the application of knowledge graphs has involved improving recommendation systems to better recommend hotels to users [Cadeddu2024Optimizing]. In the broader business context, Knowledge Graphs have been attracting increasing attention because of their inherent usefulness to aid with Retrieval Augmented Generation [edge2025localglobalgraphrag].
@@ III-A Dataset
The dataset we used is a fraction of the HotelRec [antognini2020hotelrec] dataset. Which contains roughly 50 million reviews on hotels from TripAdvisor. We used only the first 10,000 reviews in this dataset for this study. Each review in the database contains the following features:
@@
Among these 10,000 reviews were 59 different unique hotels. The average rating of all these reviews is 4.162, with a standard deviation of 1.08. The average length of a review is 133 characters, while the median review length is 94. Any reviews longer than 1000 words were removed from this chart for readability purposes. For data split, to test the validity of our results we also make use of 10-fold cross validation in addition to our 80-20% train-test split, as it has been deployed in the literature for model optimisation and de-biasing the model testing [moreau2018semantic, maldonado2017detection].
@@ III-B1 Preprocessing and Embedding
The preprocessing pipeline is implemented using standard tools from the nltk [bird2009natural] and scikit-learn [pedregosa2011scikit] libraries. The result is a set of clean, tokenized text inputs ready for feature extraction using the respective embedding techniques.
@@ III-B2 Sentiment Analysis
Optionally, VADER [hutto2014vader] sentiment analysis can be used to generate a sentiment score for the review, and use this as an additional feature when training the model.
@@ III-C1 Preprocessing and Triple Extraction
Each review is first translated to English, if necessary, using automatic language detection and the Google Translate APwe [googletrans]. Next, contractions are expanded, for example: “can’t” $\rightarrow$ “cannot”, to standardize expressions. The text is then cleaned by removing HTML tags, hyphens, apostrophes, and similar punctuation that may interfere with parsing.
@@ III-C2 Sentiment Analysis
Optionally, sentiment analysis can now be performed on the list of triples using the VADER sentiment analyzer [hutto2014vader]. This returns a sentiment score that can range from -1 to 1. This score is stored in the same csv file, in a separate column.
@@ III-C3 Graph Embedding
Once that is done, for each object and subject a node is created with a relationship created based on the predicate. Each of these nodes is connected to a node representing the review. The review node in turn is connected to a node representing the hotel. If the object/subject is in the list of amenities, it instead gets a special node type called "amenity", otherwise its node type is called "word". This is only relevant for the system visualization. Once this graph has been constructed using NetworkX [hagberg2008exploring]. It is then converted into a Neo4J [neo4j] graph. This is to be able to visualize the graph with many more nodes than NetworkX supports, and to make use of its Node2Vec feature.
@@ III-D LLM-Based Model
Our LLM-based model works by asking the LLM to predict the ratings of a set of reviews, based on the initial "training set" it is given, example-based learning [ren2025malei, romero2025manchester]. So we first give it 200, or 2000 reviews with their corresponding ratings, and then based on those it has to predict the ratings for the other 9800, or 8000 reviews we give it next. To accomplish this task the LLM is allowed to write its own code.
@@ IV-B ReviewGraph Results
For the following experiments, we exclusively use the Node2Vec embedding method, as implemented in the Neo4j Graph Data Science Library [neo4j]. The model was configured with the following parameters: 1 iteration, embedding dimension of 10, walk length of 80, return factor of 1, in-out factor of 1, and a window size of 10. The embeddings were generated for nodes with the label review, using relationships of type Any and orientation Any. No relationship weight property was applied in this setup. These values may change if mentioned.
@@ IV-E Discussion
V Visualising ReviewGraph
The graph structure of ReviewGraph demonstrates significant potential as a valuable data visualization tool for understanding customer feedback. Currently, we developed several minor features for this graph visualization, which lets the user sort the graph to show data they are specifically interested in. More features can easily be added in the future. Some of the current features are:
  • Filtering reviews by a specified score or score range (e.g., 2–4).
  • Highlighting reviews, hotels, or keywords based on highest or lowest sentiment scores.
  • Enabling users to inspect low-sentiment reviews by displaying their associated nodes, providing a high-level overview of the review’s content.
Figure 4 and Figure 5 are examples of how the visualisation currently works. Every node is interactable and it shows all the nodes it is connected to when clicked on.
Figure 4: Example of a hotel and two negative sentiment reviews connected to it. The blue circle is the hotel, the green ones are the reviews, and the yellow ones are subjects/objects extracted from the review text.
Figure 5: Example of review and all the nodes connected to it as well as the relationships between those nodes connected to it. The blue one is the hotel, the green one the review, and the yellow ones are subjects/object extracted from the review text.
VI Conclusions
@@
VI-A Limitations and Further Research
@@
Sample of Extracted Triples
@@
Complete Tables of Results
-B Train Test Split Results
@@
-C 10-Fold Cross Validation Results